3. Let Yİ ~Gamma ( -3,ß-3), Y ~Gamma( -5, ß-1), and W-2% + 6K. a) (9 pts) Find the moment generating function of W. Justify all steps b) (3 pts) Based on your result in part (a), what is the distribution of W(name and parameters)?
Answer #1

as mgf of gamma distribution Mx(t) =1/(1-\betat)\alpha

hence mgf of Y1 =My1(t)=1/(1-3t)3

therefore mgf of 2Y1 =M2y1(t) =My1(2t)=1/(1-3*2t)3 =1/(1-6t)3

similarly mgf of 6Y2=M6y2(t) =1/(1-1*6t)5 =1/(1-6t)5

hence mgf of W =Mw(t)=M2y1(t)*M6y2(t) =(1/(1-6t)3)*(1/(1-6t)5) =1/(1-6t)8    t<1/6

b)

comparing mgf of W with that of gamma distribution of Mx(t) =1/(1-\betat)\alpha

W follows gamma distribution with parmeter \alpha=8 and \beta=6